OpenStudy (anonymous):

for what values of m does the line with equation y = mx - 12 not intersect or touch the parabola with equation y = 2x^2-x-10. Please show working out & explain.

OpenStudy (anonymous):

This has got me stumped and I need to know too, lol!

OpenStudy (anonymous):

I'm gonna post this on another forum and see if I get a response.

OpenStudy (anonymous):

if you're in Calculus ... could be done easily ... not sure what math are you taking...

OpenStudy (anonymous):

I'm a total dork. I took precalc 5 years ago. So I'm embarrassed

OpenStudy (anonymous):

math, if you remember then can you post the solution or at least part of it?

OpenStudy (anonymous):

LOL... this is quadratics functions just to let u know. Hopefully u will be able to answer my question soon enough. I am also taking Maths Methods and do not know what is calculus lol.

OpenStudy (anonymous):

Well they talk about this stuff in precalc.

OpenStudy (anonymous):

So mathg8 is telling me that "I should know this" but it looks like he's not posting ANYthing.

OpenStudy (anonymous):

thinking how would i do it with algebra ...

OpenStudy (anonymous):

sorry I'm not very familiar with calculus & precalc stuff. I just know this was one of my questions I had to do on my practice test or SAC if u know what it is and ur weren't allowed to use a calculator for it, so I want the working out/solutions as to how it can be solved.

OpenStudy (anonymous):

0 would work for sure

OpenStudy (anonymous):

below the vertex

OpenStudy (anonymous):

well yeah

OpenStudy (anonymous):

I'm interested in the algebraic method of solving this. I feel completely lost on it and it's bugging me.

OpenStudy (anonymous):

do you need one value or a range of values?

OpenStudy (anonymous):

I'll bring the problem in to my old calc teacher and she'll probably beat me with an eraser!

OpenStudy (anonymous):

LOL. Well I do have the answer that I copied from the board which the teacher wrote. The teacher only wrote the answer. I can tell u the answer but I wanted to know if u could show me the working out and confirm the answer for me instead of providing u with the answer straight away. The answer is a range of values btw.

OpenStudy (anonymous):

so it is calculus

OpenStudy (anonymous):

take the derivative

OpenStudy (anonymous):

I took calc 1-3 about 5 years ago. And I got into math again just recently.

OpenStudy (anonymous):

yes Quadratic Functions, but try solving it without calculator (as I was not allowed one) and show me all the working out with it.

OpenStudy (anonymous):

y= 4x-1 ( the slope of all tangent lines to the parabola )

OpenStudy (anonymous):

the line parallel to one of the tangents has the same slope

OpenStudy (anonymous):

Jay, what answer did you get?

OpenStudy (anonymous):

m=4x-1 as you plug in values for x you'll find the values ... I'm not seeing any restrictions for x

OpenStudy (anonymous):

It's not my answer, it's the teacher's answer just to let u know. It's -3<m<5. Please help me obtain the solutions/working out, I urgently need to understand this.

OpenStudy (anonymous):

so there are some restrictions ... I was wondering and just tried -1 ...doesn't work

OpenStudy (anonymous):

So I guess you could set mx-12 = 2x^2 -x-10 and solve for m. Those will be the slopes which the line intersects I believe? So then you pick the range where it does not intersect.

OpenStudy (anonymous):

-1 is between -3 and 5 and it doesn't work ...

OpenStudy (anonymous):

I'm not too sure but this is the answer that I copied from the board.

OpenStudy (anonymous):

-5<m<3

OpenStudy (anonymous):

did you get 3 and -5 switched ?

OpenStudy (anonymous):

I'm not sure that I copied it wrong or if the teacher has the wrong answer. I do not even know how to get to the answer.

OpenStudy (anonymous):

2x^2 -x -10 =mx-12 find the determinant and make it < 0 ( 2 imaginary solutions) 2x^2-x-mx-10+12 = 0 D = (1+m)^2 -16 < 0 (1+m)^2 < 16 1+m < 4, m< 3 1+m < -4 , m<-5 -5<m<3